OpenStudy (anonymous):

The product of two consecutive positive integers are 1 more than their sum find the integers.

Parth (parthkohli):

Or, use trial and error method.

Parth (parthkohli):

It works with 2 and 3.

OpenStudy (anonymous):

consecutive is the missing part

OpenStudy (anonymous):

since they're consecutive, the equation is x(x+1) = x + (x+1) + 1

OpenStudy (anonymous):

\[x(x+1)=x+x+1+1\]

Parth (parthkohli):

@satellite73 Oops :P

Parth (parthkohli):

So this doesn't need a system.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x(x + 1) = 1 + x + x + 1 }\)

OpenStudy (anonymous):

is the answer -1,2

OpenStudy (anonymous):

but you are right you can probably do it by trial and error, although this is a quadratic so there should be two solutions \[x^2+x=2x+2\] \[x^2-x-2=0\] \[(x-2)(x+1)=0\] \[x=-1, x=2\] and so the two solutions are 2 and 3 or -1 and 0

Parth (parthkohli):

@satellite73 You missed the POSITIVE part. Oops,

OpenStudy (anonymous):

aaah so i did it does say positive, doesn't it solution is 2 and 3 as you said right at the start